# A Simple Model of Current Ramp-Up and Ramp-Down in Tokamaks

### v1

@@ I Introduction
    Did not explicitly include the changes in the flux in the central solenoid in his analysis, which is the only flux that is a directly controllable function of time.
@@
Fitzpatrick made no note of the sentence in the abstract of the submitted version of my article that gave the major implication on plasma shutdown: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.” Fitzpatrick did not attempt to determine what range of plasma current profiles are stable, and he omitted the majority of the poloidal flux produced by the plasma current, the part outside of the plasma.
Fitzpatrick’s certainty that my concerns about disruptions are unfounded relate to a statement in his September 10, 2025 email to me: “the average broadness of the current profile is easily controlled.” This statement by Fitzpatrick might make one think disruptions in tokamak would be rarely encountered and then only as a result of operator incompetence. It was not so simple in JET shutdowns Sozzi:2020 .
Section III discusses Fitzpatrick’s misrepresentation of my views. As stated by the first reviewer of my Physics of Plasmas submission Boozer:flux : “a more rigorous model was recently published in Nuclear Fusion by Fitzpatrick, prompted by the arXiv article from this author.” Fitzpatrick’s first reference was my arXiv article. His second reference was “Boozer A.H. 2025 private communication.” The “private communication” was a number of emails that also made my concern explicit—the smallness of the fractional poloidal flux change required to produce a disruptive state.
The extreme nature of the errors in science and of the misrepresentation of my views caused me to enquire, starting on November 24, 2025, whether Fitzpatrick’s paper is consistent with the scientific and ethical standards of Nuclear Fusion. On January 5, 2026 Nuclear Fusion said they had “no evidence of intentional misrepresentation,” despite the emails.
The urgency of having this Comment publicly available is explained in Section IV. Both the renewal of grants and the publication of a paper on quick assessments of issues faced by tokamak power plants and of strategies to circumvent them are dependent on the public availability of this Comment.
@@ II.1 Basic properties the poloidal flux
may seem odd, but it results in $\psi_{p}$ increasing with distance $r$ from the axis, which gives a positive safety factor. The distance from the axis is $r\equiv R_{ax}-R$, and $2\pi(R_{ax}-r)dr$ gives the area element for defining the poloidal flux that lies within a distance $r$ of the axis.
@@ II.2 Disentangling Fitzpatrick’s expressions
The first step in a disentanglement is determining what is precisely Fitzpatrick’s definition of the poloidal flux and what basic equations does it obey in the cylindrical $(r,\theta,z)$ approximation that he utilizes. His Equations (6) and (9) together with his Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the Fitzpatrick definition of the poloidal flux $\psi_{f}(r,t)$ is
@@
The poloidal magnetic flux is the flux enclosed in the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$. The time derivative of the poloidal flux satisfies exactly
@@ II.3 Fitzpatrick check for non-disruptive profiles
What Fitzpatrick did not do is determine the range of profiles of his energy diffusivity that give non-disruptive current profiles even by the Cheng et al criterion that he was using.
Showing stability for one profile of the heat diffusivity is certainly no guide for determining the danger of disruptions. It might be noted, that the general view among tokamak experts is that energy diffusivity is not the correct way to characterize energy transport. The temperature profiles are given rigidity by microturbulence. See for example Reference Holland:2021 . This rigidity is sometimes thought to ensure that a spatially constant loop voltage will give disruption-stable current profiles.
@@ II.4 Actual condition for disruption avoidance
III Fitzpatrick’s misrepresentations
Fitzpatrick made it clear in the Abstract, Introduction, and in the Summary and Discussion that he wrote his paper Fitzpatrick:2026 to correct an arXiv article of mine, his first reference. That would be fine except he was misinterpreting my article. His second reference was “Boozer A.H. 2025 private communication.”
The “private communication” was a chain of emails some even before Richard submitted his article to Nuclear Fusion on August 18. These emails politely but explicitly stated that he was misinterpreting me, that I was revising my article to remove any confusion, and what I viewed as the most important point of my article. If Fitzpatrick had just ignored these emails, that would have been irresponsible. But, Richard coupled his first and second references everywhere except in the Abstract to reinforce his claim to be correcting a false view of mine.
His misrepresentation of “private communication” is particularly clear in an August 13 email exchange because it was before his paper had been submitted to Nuclear Fusion. My email to Fitzpatrick said “Actually I didn’t really mean the expected ramp down time was 1000 sec, and it would take a lot of power to keep the plasma going. What I was trying to indicate is that a large cooling of the plasma is necessary but the profile had to remain relatively fixed while doing that ramp down. I agree my words were confusing, and they have been changed.” Fitzpatrick responded: “Would it be accurate to say that you think that the ramp-down times are longer than the SPARC and ITER people are claiming” My response: “The primary issues are the energy and magnetic flux input into the plasma that is required during a shutdown to ensure the current profile does not slip into a disruptive state more than one time in at least a thousand shut-downs. A major problem in my view, is that a 20% change in the poloidal flux distribution is sufficient to take a plasma far from a state as far as possible from a disruption into a disruption. This means a disruption can occur in approximately a fifth of the time one would expect a flattop to last.”
The perception of the first reviewer of my Physics of Plasmas submission Boozer:flux is revealing: “a more rigorous model was recently published in Nuclear Fusion by Fitzpatrick, prompted by the arXiv article from this author.” Fitzpatrick told me that he was not the reviewer in a response to a version of this Comment that I sent to him for comments. The first reviewer perceived that the whole motivation for Fitzpatrick’s paper Fitzpatrick:2026 was my arXiv article but did not consider the statement of purpose in the abstract of that article.
The last “private communication” that Fitzpatrick and I had was on 9-10 September 2025, which was before he submitted his revised version of his article, 25 September 2025. This was the version that was accepted. Since this communication illustrates our different perceptions of physics issues, it is quoted exactly below. In his article, Fitzpatrick did not ascribe to me the views that I stated, nor does he address in this email or in his paper his confusion about the definition and the evolution of the poloidal flux.
Boozer-Fitzpatrick 9-10 September emails
From: Richard Fitzpatrick rfitzp@utexas.edu Subject: Date: Re: Revised abstract & introduction to “Constraints on the magnetic field evolution in tokamak power plants” September 10, 2025 at 4:27 PM To: Allen H. Boozer ahb17@columbia.edu
Dear Allen I do not really agree that tokamak plasma are “self-organized”. When I think of the term self-organized, I naturally think of an RFP plasma. Such plasmas are subject to continuous large-amplitude m=1 and m=0 tearing modes. It is these modes that relax the plasma toward a Taylor state. However, a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central q value from falling significantly below unity. Given that q(0) is pinned to a value close to unity, and q(a) is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities. Richard
On Tue, Sep 9, 2025 at 4:29 PM Allen H. Boozer <<ahb17@columbia.edu>> wrote: The confusions on the important points in my arXiv article “Constraints on the magnetic field evolution in tokamak power plants” were convincing that the abstract and introduction had to be rewritten. Any comments are appreciated, especially ones on balance and fair quoting. Discussions by Zoom or phone are always welcome, Allen 646-319-9962
A copy of my article Boozer:flux was attached to my email and a plaintext version of its abstract was included.
IV Urgency
Two issues make the publication of this Comment urgent.
  1. 1.
    The U.S. Department of Energy rejected the renewal proposal that would have provided support for this paper on December 1, 2025 after it had been under review for twenty-three months. This was shortly after Fitzpatrick’s Nuclear Fusion paper appeared online, November 6, 2025. It is not only important that my funding for this type of work be restored, but also one other renewal proposal be funded, which has been under review for more than eight months, has not yet had a response.
    The renewal of proposals based on a continuing development of fundamentally new concepts and inventions has become difficult under the new Fusion Science and Technology Roadmap DoE:2025 . Page 7 of the Roadmap prominently displays: “Innovate and advance the science and engineering of fusion with well-defined milestones and metrics.” As interpreted, “well-defined milestones” is consistent with funding based on the development of fundamentally new concepts and inventions only when their timing can be predicted years into the future—a logical impossibility.
    Having my contributions trashed in a respected journal, even in an article filled with scientific errors and misinterpretations does not help. It is natural to assume that Nuclear Fusion, a respected journal, would not publish anything with so many problems—and would not stand behind such a paper if it were published.
  2. 2.
    The Physics of Plasmas has said that they would consider a new submission of my article Constraints on the magnetic field evolution in tokamak power plants only after the incorrectness of the charges in Richard Fitzpatrick’s paper against my article had been demonstrated.
    My article allows quick assessments of issues faced by tokamak power plants and of strategies to circumvent them. At present a number of privately funded proposals for tokamak power plants are being developed, and such assessments are of great practical importance. There is no reason that a paper filled with scientific mistakes and misrepresentations should be allowed to continue to block a publication important to power-plant proposals.

### v2

@@ I Introduction
    Did not explicitly include the changes in the flux in the central solenoid, $\psi_{sol}$, in his analysis, which is the only flux that is a directly controllable function of time.
@@
Fitzpatrick has insisted on misinterpreting my article as primarily about acceptable ramp-up and ramp-down times. The abstract of his paper states: “there is no indication that the design ramp times are infeasible, as was recently suggested in Boozer (2025 arXiv:2507.05456v1).” Somehow he came to this conclusion when even the first arXiv version, which he emphasizes rather than the version submitted to the Physics of Plasmas, stated the long decay time of the plasma current at fusion temperatures means: “A rapid cooling of the central part of the plasma seems required, but unless this is done with care, the current profile will evolve into a disruptive state.” The same arXiv version also said the restart period in a pulsed tokamak has significant freedom: “In principle, the restart period can be made arbitrarily short by clever choices of the time dependence of $\psi_{sol}(t)$ and the heating power. Of course, the periods in which substantial heating power is required must be short compared to the periods of fusion burn to have net energy to sell.” Fitzpatrick made no note of the sentence in the abstract of the submitted version of my article: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
As discussed in Section II.4, the smallness of the required change in the poloidal flux to cross the full range of stable plasma current profiles implies the current profile in the plasma is determined by the spatial constancy of the loop voltage. Actually Fitzpatrick assumed a constant loop voltage to obtain the single profile of plasma current that he discussed in his Nuclear Fusion paper, but did not explain why this was an accurate approximation.
Despite a number of emails that attempted to correct Fitzpatrick’s mistaken interpretation of my article, he not only persisted in that misinterpretation but reinforced his misinterpretation in every reference to my arXiv article, except in the abstract, by citing “private communication” as well. The only private communication that we had was the emails that tried to correct his misinterpretation.
The reason for Fitzpatrick’s persistence on misinterpreting my article can be explained by his statement of dogma in his September 10, 2025 email, part of his “private communication” reference: “a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central $q$ value from falling significantly below unity. Given that $q(0)$ is pinned to a value close to unity, and $q(a)$ is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities.”
Fitzpatrick’s dogma seems to imply that disruptions could be easily avoided by a competent machine operator. It was not so simple in JET shutdowns Sozzi:2020 .
Although I have never said that tokamaks have “an inevitable drift to a disruptive state,” ensuring tokamaks have a sufficiently low disruption rate for the economic feasibility of power plants is non-trivial and requires improved understanding.
A similar dogma to the one that prevented Fitzpatrick from acknowledging my views seems to affect the judgement of a significant part of the tokamak community and allows them to overlook even obvious mistakes in science. To correct this, the scientific evidence that does not conform with Fitzpatrick’s statement of dogma needs to be given.
A Physics of Plasmas reviewer of my article said Fitzpatrick’s paper is “a more rigorous model.” A leader in the tokamak program stated that Fitzpatrick’s paper: “refutes your arguments that this (disruptions) inherently dooms tokamaks. This reason goes well beyond Richard’s paper, which attacks the problem from a fundamental and transparent way in a simple model. This is because some of the most sophisticated and benchmarked numerical simulations we have for tokamaks are time-dependent stability/equilibrium codes that tackle the inductive problem. These have been developed over decades of development and testing against thousands of tokamaks discharges on an extremely large range of device sizes, plasma temperatures and engineering design.” Similar comments have come from a number of persons.
It is surprising that anyone believes the quality of the science underlying the design of tokamak power plants is at such a low level that Fitzpatrick’s paper adds important information about disruption avoidance. The glowing descriptions of “sophisticated and benchmarked numerical simulations we have for tokamaks” is an appropriate statement to make to a potential private funder but is not a scientific response to a scientific study into the requirements for avoiding disruptions.
In addition to theory and simulations, empirical studies advance understanding of disruptions and their avoidance. Studies based on the DCAF code developed by Steve Sabbagh’s group have been used to determine Sabbagh:2024 the empirical dependence of tokamak disruptions on plasma quantities such as the internal inductance $\ell_{i}$ and edge safety factor $q_{95}$, Figure 2, The results for MAST-U, which has highly non-circular magnetic surfaces, give preliminary evidence that an internal inductance that is either too low or too high gives disruptive plasmas with a stable region between. An empirical study that is related to that of Figure 2 is given in Figure 3, which is a study of disruptions in JET with an ITER-like wall, JET-dis:2020 . That figure shows that internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ are not sufficient to define a stable region. Are the “sophisticated and benchmarked numerical simulations we have for tokamaks” adequate for explaining Figures 2 and 3?
Figure 2: MHD-caused disruptions are illustrated in an internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ diagram with separate symbols for KSTAR and NSTX-U as well as for two operational years of MAST-U. Two branches of DCS trigger points can be recognized in both MAST-U experimental years. The two branches (visualized with green lines) are separated along the $\ell_{i}$ axis into two regions, with the internal inductance being roughly restricted to $1.5\lesssim\ell_{i}\lesssim 1.7$ in branch 1 and $0.8\lesssim\ell_{i}\lesssim 1.2$ in branch 2 . The region between the branches is free of MHD-caused disruptions for $q_{95}\gtrsim 6.3.$ This is Figure 7 in the paper V. Zamkovska, S.A. Sabbagh, M. Tobin, et al, Nucl. Fusion 64, 066030 (2024).
Figure 3: Pre-disruptive parameters in JET with an ITER-like wall are shown in a $\ell_{i}-q_{95}$ stability diagram. This was Figure 13 in Gerasimov et al, Nucl. Fusion 60, 066028 (2020). As the authors note: “It may be expected that a disruption free space may be defined in the $\ell_{i}-q_{95}$ empirical stability diagram, assuming that plasma current profiles tend to maintain itself inside the permissible values. In reality, the JET-ILW pre-disruptive plasma equilibrium parameters create a diffused cloud on the $\ell_{i}-q_{95}$ stability diagram without room for non-disruptive plasmas.”
Ensuring disruptions are sufficiently rare for the feasibility of tokamak power plants would seem to require identifying the externally controllable parameters that can be localized in a way that makes the plasma safe against disruptions. This requirement is more stringent than in existing tokamaks or in JET because of the limitations on diagnostics and actuators in power plants. Generative AI using data from both experiments and simulations could help in this identification.
A major determinant of disruptivity is the profile of the net plasma current $I(\psi_{t})$ enclosed by a magnetic surface that also encloses a toroidal magnetic flux $\psi_{t}$. The derivative $I^{\prime}(\psi_{t})\equiv dI/d\psi_{t})$ is given by the divergence-free part of $\vec{j}_{||}=(j_{||}/B)\vec{B}$, which is the net plasma current, and is determined by the spatial constancy of the loop voltage. Despite the dogma expressed by Fitzpatrick and apparently accepted by many in the tokamak community, the naturally arising profiles of $I(\psi_{t})$ during startup, flattop, and shutdown have many uncertainties. Figures 2 and 3 show that actual tokamak plasmas have many values of $\ell_{i}$, which quantifies the peakedness of the current profile, for a given edge safety factor. There is not just the one current profile that Fitzpatrick focused on.
What theory can do and what I am doing is identifying controllable features in $I(\psi_{t})$ that determine the internal inductance $\ell_{i}$. When the central safety factor $q_{0}$ is held to unity by sawteeth, the derivative $I^{\prime}(0)$ is defined. The determination of the current density $I^{\prime}(\psi_{t})$ by the loop voltage implies $I^{\prime}(\psi_{t})\geq 0$ without negative external current drive. Both a positive $I^{\prime}$ near the edge, as in a H-mode plasma, and triangular shaping of the magnetic surfaces reduce $\ell_{i}$ for a fixed edge to central $q$-factor ratio with an otherwise fixed $I(\psi_{t})$. It would be interesting to determine if extending the empirical data used to construct Figures 2 and 3 would be much more determinate using a four than in a two parameter fit.
The acceptance of a Fitzpatrick-like dogma apparently led to strong negative reactions to two statements that were in the various versions of my arXiv article: The first statement was: “Diversity of concepts is important for the development and optimization of fusion. Nevertheless, an intelligent allocation of resources among the various concepts requires theoretical and computational assessments of the risks, time, and costs of fusion development based on these concepts.” In principle, it would seem anyone who believes in the importance of the rapid development of fusion would agree. The second statement was: “The purpose of this paper is not to remove tokamaks from consideration but rather to determine on what issues research should be focused for tokamaks to be a credible alternative to stellarators. For this, the obvious strategy is to determine what actuators are available to control the things that must be controlled and what plasma data could be available to guide these actuators.” This statement seems controversial only when one believes the external-current determination of stellarator plasmas is not fundamentally different in disruption and other plasma control issues when compared to the effects of internal plasma currents in tokamaks.
Anyone responsible for managing the development of fusion power plants would do well to ponder Admiral Rickover’s well-known quotes based on his success in the launching the first nuclear submarine about fifteen years after the splitting of the uranium nucleus in Berlin. An important quote is: “It is a human inclination to hope things will work out, despite evidence or doubt to the contrary. A successful manager must resist this temptation.”
@@ II.1 Basic properties the poloidal flux
with $Z=0$, may seem odd, but it results in $\psi_{p}$ increasing with distance $r$ from the magnetic axis, which gives a positive safety factor. The distance from the axis is $r\equiv R_{ax}-R$, and $2\pi(R_{ax}-r)dr$ gives the area element for defining the poloidal flux that lies within a distance $r$ of the axis.
@@ II.2 Disentangling Fitzpatrick’s expressions
The first step in a disentanglement is determining what is precisely Fitzpatrick’s definition of the poloidal flux and what basic equations it obeys in the cylindrical $(r,\theta,z)$ approximation that he utilizes. His Equations (6) and (9) together with his Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the Fitzpatrick definition of the poloidal flux $\psi_{f}(r,t)$ is
@@
The poloidal magnetic flux is the flux enclosed in the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$.
The time derivative of the poloidal flux outside a magnetic surface that encloses toroidal flux $\psi_{t}$ satisfies exactly
@@ II.3 Fitzpatrick check for non-disruptive profiles
What Fitzpatrick did not do is determine the range of profiles of his energy diffusivity that give non-disruptive current profiles even by the Cheng et al criterion that he was using. Showing stability for one profile of the heat diffusivity is certainly no guide for determining the danger of disruptions.
In his response to an earlier version of this Comment, Fitzpatrick stated Fitzpatrick:2026A : “The transport of electron energy in tokamak plasma has been the subject of extensive research for the last 50 years. There is overwhelming evidence that such transport is diffusive in nature.”
Despite the dogmatism of Fitzpatrick’s statement, the general view among tokamak experts is that energy diffusivity is not always the correct way to characterize energy transport. The temperature profiles can be given rigidity by microturbulence, Holland:2021 ; Turb-profiles:2025 . The study by N.T. Howard et al of microturbulent transport in ITER Turb-profiles:2025 has sufficient information to calculate the loop voltage and determine its spatial constancy. This would determine the consistency of their assumed current profile $I(\psi_{t})$. Much work remains to establish the consistency of current profiles with both thermal transport and stability against disruptions in full ramp-up, flattop, and ramp-down scenarios even as the nuclear power production turns on and off.